OpenStudy (anonymous):

Can someone PLEASE give me a 7th grade algebra equation, and show me how to solve it?

OpenStudy (amistre64):

7x + 10 = 59

OpenStudy (anonymous):

2x + 1 = 3 undo +1 by -1 from both sides to get balanced 2x=2 divide both sides by 2 x=1

OpenStudy (anonymous):

So, you would have to go through the order of operations, and start with the... ugh

OpenStudy (anonymous):

i'm so stupid >:(

OpenStudy (anonymous):

inverse operation + and -, x and /

OpenStudy (amistre64):

we can start simpler: how about: x+3 = 5

OpenStudy (anonymous):

that would be 2

OpenStudy (amistre64):

you are correct, but why is it 2?

OpenStudy (anonymous):

because 3- 5=2

OpenStudy (anonymous):

always undo addition or subtraction first then multiplication or division...you got it!

OpenStudy (anonymous):

:D so, can you give me a harder one?

OpenStudy (amistre64):

almost correct :) 5 - 3 = 2

OpenStudy (anonymous):

5-3 = 2

OpenStudy (anonymous):

wait what did i say?

OpenStudy (amistre64):

3 - 5 = -2

OpenStudy (anonymous):

oops thats what i meant... that would be a negative right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

very good!

OpenStudy (amistre64):

but lets step thru this easy one so you can know what is going on underneath it all. It does no good to simply go ..thats an easy one! and give the answer right?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

P.E.M.D.A.S.

OpenStudy (amistre64):

x + 3 = 5 that equal sign means that we want to keep things in balance, have you ever played on a seesaw?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

what happens when the kids on each end weigh the same?

OpenStudy (anonymous):

one will have to push off to make them move, or else it will be balanced. and that would still be 2

OpenStudy (amistre64):

correct :) now, we want to keep our equation balanced at ALL TIMES!! otherwise it becomes a lie...and lies are bad :)

OpenStudy (anonymous):

yes... haha

OpenStudy (amistre64):

if we want to solve for "x" we need to get it all by itself so we can ask it some personal stuff.... we need to get rid of all that stuff that is attached to it right?

OpenStudy (anonymous):

yeah... so you would have to follow order of operations?

OpenStudy (amistre64):

order of operations is good to know, and we will get to that sometime.. but lets take this step by step so we understand the basics that every other problem you come across will face ok?

OpenStudy (anonymous):

ok. i'm sorry i suck at math :/

OpenStudy (anonymous):

those are the basics just remember to use the order of operations backwards ; )

OpenStudy (amistre64):

so lets look at the (x+3). its got extra stuff; it has a (+3) sticking to it right? If we could get this to be more like (x+0) that would be great right?

OpenStudy (anonymous):

yes i think?

OpenStudy (amistre64):

does x+0 = x?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then we want x+0 :) how do we get rid of (+3) and turn it into "0"??

OpenStudy (anonymous):

ummmmm.... put it in parenthises?

OpenStudy (anonymous):

NO

OpenStudy (amistre64):

i wish lol

OpenStudy (anonymous):

WE WOULD SUBTRACT

OpenStudy (amistre64):

thats correct, we would take away 3. So lets do that here: x+3-3 = 5-3 because if we take away from one side we gotta take away from the other right?

OpenStudy (anonymous):

yeah. so x would equal 2?

OpenStudy (amistre64):

it will....but lets see why. we get x+0 = 2 from this. And we all no that zero is pretty useless...not!! x+0 = x, and thats important to know :) so x = 2

OpenStudy (anonymous):

ahhh i understand.

OpenStudy (anonymous):

wait where did you get the 5?

OpenStudy (anonymous):

oh no nevermind.

OpenStudy (amistre64):

I dug around in the recesses of my brain and pulled it out to create a 7th grade algebra question for you lol

OpenStudy (anonymous):

:D

OpenStudy (amistre64):

tell me what I do with this one now: t - 4 = 3

OpenStudy (anonymous):

turn the 4 into 0? i think i know the answer to this though

OpenStudy (anonymous):

1?

OpenStudy (amistre64):

lets turn the 4 into how? we have t - 4 .... what doe we do to turn it into t+0 ??

OpenStudy (anonymous):

O.O

OpenStudy (amistre64):

4 into 0 how...

OpenStudy (anonymous):

oh subtract

OpenStudy (amistre64):

lets try that and see if it works ok? t - 4 - 4 = 3 t - 8 = 3 .... that didnt do us much good did it? lets go back to: t - 4 = 3

OpenStudy (anonymous):

ok. so would we change it to 4-3=t?

OpenStudy (amistre64):

stay with me now :) do you see that t is attached to a (-4)? t-4

OpenStudy (anonymous):

¿:ı

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a negative 4?

OpenStudy (amistre64):

good, yes :)

OpenStudy (amistre64):

how do we get rid of a -4?

OpenStudy (anonymous):

add 4

OpenStudy (amistre64):

exactly :)

OpenStudy (amistre64):

lets try it out: t -4+4 = 3+4 t-0 = 7 t = 7

OpenStudy (anonymous):

i think i understand that.. but why wouldn't the 3 be a negative if its... wait i'm confusing myself....

OpenStudy (amistre64):

we add 4 to EACH side right? t -4+4 = 3+4 what does 3+4 =?

OpenStudy (anonymous):

7....

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

i understand

OpenStudy (amistre64):

good :)

OpenStudy (amistre64):

wanna try another one?

OpenStudy (anonymous):

thanks you for doing this for me. and yes please

OpenStudy (anonymous):

2x+8=16

OpenStudy (anonymous):

you have to do the oppsite \

OpenStudy (anonymous):

¿:ı

OpenStudy (amistre64):

youre welcome: lets try something related, but different now. we will get to that one soon enough

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

5x = 15

OpenStudy (anonymous):

sounds good

OpenStudy (anonymous):

-8 both sides

OpenStudy (anonymous):

2x+8=16

OpenStudy (anonymous):

ummm 5x would be like, 5• 5?

OpenStudy (amistre64):

it would be the same concept as 5*5 right.

OpenStudy (anonymous):

so... would you be solving for x in that equation?

OpenStudy (anonymous):

2x+8=16 -8 -8 2x=8 2x/2=8/2 x=4

OpenStudy (amistre64):

yes. if we add to get rid of a subtraction. And we Subtract to get rid of an addition; what do youthink we might do to get rid og a multiplier?

OpenStudy (anonymous):

oh xcrap.. divide?

OpenStudy (anonymous):

i hpoe this helps

OpenStudy (amistre64):

divide is correct:)

OpenStudy (anonymous):

2x+8=16 -8 -8 2x=8 2x/2=8/2 x=4

OpenStudy (anonymous):

that one

OpenStudy (anonymous):

i suck at dividing :D thanks for being patiant

OpenStudy (anonymous):

do you get my eample

OpenStudy (amistre64):

we dont have to know alot about division for this really; we only have to know that any number divided by itself equals 1